cashews sell for $8 a pound, and peanuts sell for $3 a pound. A store want to make 20 pounds of mixed nuts which they will sell for $6 a pound. How many pounds of cashews and how many poiunds of peanuts should they use?
Answer by ptaylor(1894)

Let x=amount of cashews needed
Then 20-x=amount of peanuts needed
Now we know that the value of the nuts before they are mixed together(8x +3(20-x)) has to equal the value of the nuts after they are mixed together(6*20). So our equation to solve is:
8x+3(20-x)=6*20 get rid of parens and simplify
8x+60-3x=120 subtract 60 from each side
8x+60-60-3x=120-60 collect like terms
5x=60 divide each side by 5
x=12 lb------------------------------amount of cashews needed
20-x=20-12=8 lb------------------------------amount of peanuts needed
CK
8*12+3*8=6*20
96+24=120
120=120
Hope this helps---ptaylor